The cardinality The cardinality of the set A, n(A) = 50 and the cardinality of set B, n(B) = 50. These sets have 3 elements in common.
a) What is the cardinality of the set: AB?
b) What is the cardinality of the set: AB?
c) What is the cardinality of the set: [AB]A
d) What is the cardinality of the Cartesian product: A X B

Answer by robertb(5830)

The cardinality of A intersection B = 3.
The cardinality of A U B is 50 + 50 - 3 = 97.
The cardinality of (A intersection B) U A = 50, or the cardinality of (A U B)intersection A = 50, whichever the case is (the symbols are not appearing on the screen!).
The cardinality of A x B = 50*50 = 2,500.